The vertex of the parabola below is at the point -2,-1 and the point -1,1 is on the parabola. write the equation
1
Expert's answer
2013-03-01T03:53:53-0500
let y=a*(x-k)^2+b
we know that ( -2, -1 ) is the vertex and (-1,1) lies on parabola therefore these points should turn equation into right equality. So y=a*(x+2)^2-1 and from (-1,1) we get
1=a*(-1+2)^2 -1
a=2
so the equation is the following y=2(x-2)^2-1